OpenStudy (anonymous):

Is the line through points P(–7, 2) and Q(7, –6) parallel to the line through points R(–1, –5) and S(–3, –1)? Explain. A) No; one line has zero slope, the other has no slope. B) Yes; the lines are both vertical. C) Yes; the lines have equal slopes. D)No; the lines have unequal slopes.

OpenStudy (aroub):

The only way to find out is to find the slope.

OpenStudy (aroub):

P(–7, 2) and Q(7, –6) => -6-2/7+7 -8/14 PQ=- 4/7 R(–1, –5) and S(–3, –1) => -1+5/-3+1 4/-2 -2 RS=-2 D) No; the lines have unequal slopes

OpenStudy (aroub):

Parallel= same slope but here they do not have equal slopes so they're not parallel.